### v1

# Lrq-Fact: LLM-Generated Relevant Questions for Multimodal Fact-Checking

Abstract
Human fact-checkers have specialized domain knowledge that allows them to formulate precise questions to verify information accuracy. However, this expert-driven approach is labor-intensive and is not scalable, especially when dealing with complex multimodal misinformation. In this paper, we propose a fully-automated framework, Lrq-Fact, for multimodal fact-checking. First, the framework leverages Vision-Language Models (VLMs) and Large Language Models (LLMs) to generate comprehensive questions and answers for probing multimodal content. Next, a rule-based decision-maker module evaluates both the original content and the generated questions and answers to assess the overall veracity. Extensive experiments on two benchmarks show that Lrq-Fact achieves promising performance for multimodal misinformation detetcion. Moreover, we evaluate its generalizability across different model backbones, offering valuable insights for further refinement.
1 Introduction
Fact-checkers are experts with sufficient domain knowledge in their respective fields, enabling them to ask relevant and targeted questions to verify the factuality of information [72, 73]. This expertise allows them to assess the veracity of claims from multiple perspectives by formulating precise questions that probe both textual and visual content [55]. However, this manual approach is time-consuming, labor-intensive, and lacks scalability, especially given the increasing complexity and volume of multimodal misinformation. As misinformation continues to evolve, relying solely on manual efforts is no longer practical [8, 25, 27].
Figure 1: The proposed framework, Lrq-Fact, draws on insights from manual fact-checking process.
In response to this challenge, numerous automatic misinformation detection systems have recently emerged to address the problem [54]. While these systems provide valuable insights, most of them focus on textual misinformation only [81, 29, 8, 62]. This text-only detection approach may overlook critical visual cues. Few works [86, 22, 59] for multimodal scenarios often analyze these modalities in isolation, and then combine predictions from text and images data, thus leading to overlooking of fine-grained inconsistencies or manipulation patterns, such as minor facial expressions [37].
To overcome these limitations, we propose LLM-generated Relevant Questions for multimodal FACT-checking (Lrq-Fact), a framework designed to jointly consider the intricate inter-connections between textual and visual content. Lrq-Fact is inspired by the process of human fact-checkers, which can be abstracted into investigating two types of questions: those focused on image content and those focused on textual content [75]. Image-focused questions enable the system to carefully examine the visual content within the context of the article, ensuring that the image accurately reflects critical details such as people, objects, or events mentioned in the text. This step is crucial, as visual elements can be misleading or manipulated, even when the accompanying text appears accurate. Meanwhile, text-specific questions are used to verify the factual accuracy and plausibility of the claims made in the article itself.
Lrq-Fact incorporates the power of existing Vision-Language Models (VLMs) and Large Language Models (LLMs) to make the process of asking fact-verifying questions more scalable and efficient. LLMs generate precise and contextually relevant questions aimed at scrutinizing the textual content, while VLMs analyze visual content to answer image-focused queries. By integrating these models in a unified framework, Lrq-Fact addresses the limitations of isolated text or image analysis, enabling deeper detection of inconsistencies across modalities. In addition to generating separate questions and answers for image and text, our framework uses a rule-based decision-maker, where LLMs examine the insights from both modalities to determine the overall veracity of the content. This approach automates the fact-checking process and provides explainable rationales for the decisions, enhancing transparency and trust in the detection outcomes.
We evaluate the performance of our framework through extensive experiments on two benchmarks for multimodal misinformation detection. The results show that our method significantly improves detection accuracy and the quality of the rationales provided, compared to existing baselines. Additionally, we assess the generalizability of our framework across different LLM backbones, offering valuable insights for further optimization. Ovoerall, the key contributions of this work are summarized as follows:
  • We propose a novel framework, Lrq-Fact, to address the scalability issue in multimodal fact-checking by automatically generating multi-aspect questions.
  • Lrq-Fact incorporates a rule-based decision-maker that combines insights from both modalities, providing comprehensive rationales to examine overall veracity.
  • Extensive experiments on two benchmarks show improved detection accuracy and rationale quality compared to baselines. We also evaluate the generalizability of our framework across different LLM backbones.
2 Related Work
2.1 Fact-Checking
Fact-checking is a critical task in combating misinformation, traditionally performed by human experts who generate questions and answers to verify the veracity of claims [18]. This process involves evaluating statements, comparing them with credible sources, and identifying inconsistencies [19]. However, manual fact-checking is labor-intensive and does not scale well with the growing online content [72]. Automated fact-checking systems have emerged to address this issue, with early approaches focusing primarily on textual claims [20].
Several works have explored various methods for automating fact-checking by leveraging machine learning and natural language processing (NLP) [50, 26, 31]. For example, one work expanded fact-checking capabilities by integrating context retrieval with large-scale databases to verify the factuality of statements [51].
However, while text-based fact-checking has made significant strides, existing approaches often fall short when dealing with multimodal content—such as the combination of images and text commonly found on social media platforms [79, 10]. This limitation has led to the need for advanced fact-checking systems that analyze both visual and textual elements.
2.2 Multimodal Misinformation
Misinformation has become a critical issue that affects various domains, from politics to public health [24, 38, 14]. Misinformation involves spreading false or misleading information, and can exist in different modalities—such as text, images, audio, and video—making it increasingly complex to detect [70, 60, 81].
Several datasets have been developed to evaluate text-based misinformation detection. For example, the FEVER dataset [68] consists of manipulated Wikipedia sentences annotated for veracity. Similarly, FakeNewsNet [63] collects misinformation from fact-checking websites and social media platforms. However, these datasets are limited to textual modality. Some recent datasets have focused on multimodal misinformation [23, 58]. The Fakeddit dataset [49] includes Photoshop-edited images from Reddit. DGM4 [61] introduces both global and local manipulations in images to alter sentiment and semantics, providing a comprehensive evaluation of multimodal misinformation. MMFakeBench [45] is designed to assess mixed-source multimodal misinformation. By including both synthetic and naturally occurring misinformation, MMFakeBench provides a valuable resource for evaluating modern VLMs and multimodal misinformation detection frameworks. These benchmarks highlight the need for detecting both textual and visual veracity distortions, as well as cross-modal consistency, which involves verifying if the visual and textual content align correctly.
In terms of detection, recent approaches aim to evaluate the consistency between different modalities. These techniques involve learning joint multimodal representations, where both visual and textual features are fused to detect inconsistencies or manipulations. Approaches like multimodal feature representation learning [1] and image captioning models [87] attempt to capture this alignment. However, these models often suffer from a lack of interpretability, making it difficult to interpret their decisions. To address this limitation, interpretable misinformation detection has gained attention [43, 15].
2.3 LLMs and VLMs
LLMs and VLMs have revolutionized the field of multimodal misinformation detection. LLMs, such as GPT-3 [7], GPT-4 [3], and LLaMA 3.1 [71], are designed to perform a wide range of language tasks, including summarization [16, 85], knowledge discovery [4, 32, 33, 66, 80, 82], question answering [69, 84] and various applications [83, 36, 12, 28]. These models are increasingly being used in fact-checking due to their ability to analyze textual claims. Recently, GPT-4o and LLaMA 3.1, in particular, have shown significant advancements in their ability to handle diverse, real-world misinformation by offering context-aware question generation and deeper understanding of semantic structures in texts [71, 5, 48, 3].
Meanwhile, VLMs, such as CLIP [57], ViLBERT [47], and Paligemma [6], have demonstrated their ability to bridge the gap between image and text by learning representations that align both modalities. These models are good at aligning fine-grained features between text and images, making them particularly useful for tasks such as verifying if an image’s content is consistent with its corresponding caption. By leveraging the capability of these VLMs in multimodal fact-checking, systems can detect subtle manipulations between visual and textual content [6].
In this work, we integrate LLMs and VLMs for multimodal misinformation detection. Recent studies have demonstrated the potential of leveraging their ability to generate diverse QA pairs based on both text and visual content [45, 64]. Inspired by this, Lrq-Fact combines LLMs like GPT-4o and LLaMA 3.1 for language understanding with VLMs like Paligemma for image analysis.
3 Methodology
In this section, we detail each module Lrq-Fact. As shown in Figure 2, Lrq-Fact integrates LLMs and VLMs to automatically generate relevant questions and answers (QAs) for both visual and textual content. It consists of four modules: (1) Image Description Generation, (2) Image-Focused QAs Generation, (3) Text-Focused QAs Generation, and (4) Rule-Based Decision-Maker. The prompt templates we designed for VLMs and LLMs are provided in Appendix (see Figure 8, 9, 10, 11, 12, and 13).
Figure 2: The overview pipeline of our Lrq-Fact framework consists of four key components: (a) Image Description, which provides detailed contextual descriptions of the image; (b) Image-Focused QAs, aimed at assessing the accuracy of the visual content; and (c) Text-Focused QAs, which detect textual inaccuracies, contradictions, or unsupported claims. Finally, all the gathered information is synthesized in (d) the Final Judgment & Explanation module, where a role-based decision-maker generates both the prediction results and comprehensive explanations.
3.1 Image Description Generation
In the first stage of our framework, we generate a detailed description of the image to assess consistency between the visual and textual content within our rule-based decision-maker module. This step is essential for identifying misinformation, as discrepancies between the image and the accompanying text are often a key source of misleading information. To achieve this, the VLM generates a comprehensive summary of the image, capturing key elements and describing the overall context of the scene depicted.
Formally, given the image $img_{i}$, we prompt the VLM to generate its corresponding description $des_{i}$:
(3.1) $des_{i}={\rm VLM}(img_{i}).$
3.2 Image-Focused QAs Generation
In the second stage of our framework, we focus on generating image-focused questions to verify the visual content related to the news article. Leveraging the capabilities of LLMs, we formulate queries targeting key aspects of the image. These questions are grounded in the context provided by the news article and are designed to verify the alignment of the image with the claims made in the text. We achieve this by employing in-context learning, where the LLM is provided with example questions that explore critical aspects of the image, such as objects, individuals, settings, and interactions. By including these examples in the prompt, the LLM gains a clearer understanding of the types of questions required, enabling it to generate new, relevant questions tailored to the context of the article  [5]. Intuitively, this approach ensures that the generated questions comprehensively address the accuracy, consistency, and authenticity of the image.
To generate appropriate questions, we also create a well-built prompt that guides the LLM to focus on extracting relevant information from the image. The prompt aims to instruct the LLM to create high-quality, targeted questions that help fact-checkers or automated systems determine whether the visual content supports, contradicts, or diverges from the claims in the news article. These questions probe specific details such as interactions between people, the setting, actions depicted, symbols, and the presence or absence of individuals or objects mentioned in the text. For instance, consider an article that describes a sporting event: “An umpire is in jump action while the pitcher launches the ball.” In this scenario, the LLM might generate the following questions to explore the visual evidence:
  • What sport is being played in the image?
  • Is the umpire in the image wearing the appropriate attire for the sport mentioned in the news article?
  • What is the action of the pitcher in the image?
  • Are there any visible spectators in the image?
  • Is the image likely to have been generated by AI, or is it a real photograph?
Intuitively, these questions help verify whether the image accurately represents the elements described in the article, ensure its details align with the expected aspects of the event, and examine the image for potential signs of manipulation, such as determining if it was AI-generated or shows signs of tampering.
Once the questions are generated, the VLM answers them by extracting critical visual details from the image, identifying key elements such as individuals, objects, the environment, and visual cues like the presence of spectators or the setting in response to the questions. This information helps assess the accuracy of the image content in relation to the article’s claims. The abovementioned QAs generation can be expressed as:
(3.2) ${ques_{i}^{v_{1}},...ques_{i}^{v_{m}}}={\rm LLM}(text_{i}),$
(3.3) $ans_{i}^{v_{j}}={\rm VLM}(img_{i},ques_{i}^{v_{j}}),\quad 0<j<m.$
3.3 Text-Focused QAs Generation
In the third stage of our framework, we focus on verifying the textual content of the news article by generating questions that challenge the factual claims presented in the text. Leveraging the capabilities of LLMs, these questions target specific details, such as dates, names, events, and locations, to critically assess the veracity of the article. This mimics the process human fact-checkers would follow, ensuring that the text is thoroughly examined for potential inaccuracies or unsupported claims. We achieve this by employing in-context learning, where the LLM is provided with example questions that explore critical aspects of the news article, such as factual claims, timelines, or significant events. For example, if the article describes a sporting event involving an umpire and a pitcher, the LLM will generate the following types of questions to examine the plausibility of these claims:
  • Is it common for an umpire to be in a “jump action” during a pitch, and does this align with standard umpiring practices in baseball?
  • Does the description of the pitcher “launching the ball” accurately reflect the typical terminology and actions used in baseball pitching?
  • Are there any known instances or rules in baseball that would require an umpire to jump while a pitch is being thrown?
These questions are designed to probe the accuracy of the article’s claims by evaluating them against known facts or commonly accepted standards. By analyzing the responses generated for these questions, the framework can identify potential discrepancies, factual errors, or unsupported statements, ensuring that the article does not spread misinformation.
Once the questions are generated, a single LLM is used to answer them based on its extensive knowledge:
(3.4) ${ques_{i}^{t_{1}},...ques_{i}^{t_{n}}}={\rm LLM}(text_{i}),$
(3.5) $ans_{i}^{t_{j}}={\rm LLM}(des_{i},ques_{i}^{t_{j}}),\quad 0<j<n.$
Intuitively, LLMs hold a vast repository of information, far beyond what any individual can retain, and provide access to this knowledge more intuitively and efficiently than traditional search engines [75]. This makes them highly effective for fact-checking, often supplying the extra factual knowledge needed to verify claims [50]. We hypothesize that LLMs, trained on large datasets, can serve as reliable substitutes for external evidence retrieval, performing fact-checking based solely on their parametric knowledge. This allows the LLM to provide relatively reliable answers to the generated questions, assessing the truthfulness of claims in the article and identifying textual inaccuracies, contradictions, or unsupported statements.
3.4 Rule-Based Decision-Maker
In the final stage of our framework, the decision-maker integrates insights from both the textual and visual analyses to make a conclusive determination about the authenticity of the content. This stage is crucial for verifying the consistency between the image and text, and for identifying any misinformation. The decision-making process is guided by a rule-based module, designed to mimic the judgment process similar to the human fact-checkers. The LLM combines the information gathered from image and text analysis to evaluate potential inconsistencies or factual inaccuracies. The LLM will compare the image description generated by the VLM with the claims made in the news article. For example, if the article claims that a peaceful protest occurred but the image description shows signs of violence, this discrepancy is flagged. Similarly, if the article mentions a political figure being present but the image does not include this individual, the decision-maker identifies this as a significant inconsistency. This cross-modal evaluation ensures that the image and text together do not create a misleading or deceptive narrative.
The rule-based module follows a systematic process:
  • Step 1: It checks whether any of the factual claims in the text can be objectively refuted. If there is credible evidence contradicting a claim, the decision-maker labels it as “Textual Veracity Distortion”.
  • Step 2: If the text is not refuted, it evaluates the image for discrepancies. If the image contains elements that contradict facts or appear manipulated, the label “Visual Veracity Distortion” is applied.
  • Step 3: If both the text and image are individually accurate, it assesses whether they are consistent with one another. A mismatch between the text and image, such as a beach being described while the image shows a city, results in a “Mismatch” label.
  • Step 4: If none of the above issues are found, the content is labeled as “Real”.
Moreover, by providing additional guidelines ensure that the process is thorough:
  • Overall Consistency: It evaluates whether the main ideas, actions, and elements in the article align with the image description. Minor discrepancies are acceptable unless they change the overall meaning.
  • Visual Cues: Details such as body language, facial expressions, and environmental features are examined to ensure alignment with the article’s description.
  • Unrealistic Elements: It checks for AI-generated or unnatural elements in the image. Such details may indicate the image is fabricated, although the decision-maker weighs these findings based on their relevance to the article’s core narrative.
  • Final Judgment: It makes a judgment based on the overall consistency, details, and realism of both the image and the article. The decision is based on whether the story is coherent and plausible, or if contradictions and discrepancies raise doubts. Importantly, it does not over-rely on suggestions of manipulation from generated answers but makes a balanced decision based on the bigger picture.
  • Select the Most Relevant Label: After analysis, it assigns one of the following labels:
    • Textual Veracity Distortion: The text contains false or misleading claims.
    • Visual Veracity Distortion: The image description includes fabricated or unrealistic elements.
    • Mismatch: The text and image depict unrelated events or contexts.
    • Real: The text and image are consistent, without contradictions or unrealistic details.
  • Provide a Detailed Explanation: It outlines key factors influencing its judgment, focusing on text-image consistency and referencing relevant questions and answers. If generated answers hint at manipulation, the decision-maker acknowledges this but ensures it does not dominate the decision if other factors play a more significant role.
Besides providing the final judgment, the decision-maker module also provides a clear and detailed rationale explaining why the content was classified as real or false. For instance, if misinformation is detected, the decision-maker will clarify whether the decision was influenced by issues in the visual content, inaccuracies in the text, or inconsistencies between the two, based on the generated QAs. This transparent reasoning enhances trustworthiness in the fact-checking process by offering insights into how the conclusions were reached.
For each piece of analyzed content, the decision-maker provides: ❶ a final judgment $j_{i}$, such as “Real” or “Fake”, ❷ a label $l_{i}$ identifying the specific issue: “Textual Veracity Distortion”, “Visual Veracity Distortion”, or “Mismatch” and ❸ a detailed explanation of the decision-making process $e_{i}$. This rule-based approach ensures that the framework provides clear, evidence-based conclusions, carefully weighing the alignment between textual and visual information to detect misinformation. The abovementioned process can be formulated as:
(3.6) $qa_{i}^{v}=\oplus_{j=1}^{m}[ques_{i}^{v_{j}},ans_{i}^{v_{j}}],$
(3.7) $qa_{i}^{t}=\oplus_{j=1}^{n}[ques_{i}^{t_{j}},ans_{i}^{t_{j}}],$
(3.8) $j_{i},l_{i},e_{i}={\rm LLM}(text_{i},img_{i},des_{i},qa_{i}^{v},qa_{i}^{t}),$
where $\oplus$ represents the concatenation operation.
4 Experiments
In this section, we conduct extensive experiments on two benchmarks to show that our framework improves detection accuracy and rationale quality for multimodal misinformation compared to existing approaches. Additionally, we assess the framework’s generalizability across various model backbones and evaluate the robustness of different prompting strategies, providing key insights for further optimization and refinement.
4.1 Datasets
We used two multimodal misinformation detection benchmarks to evaluate the performance of our framework.
MMFakeBench: MMFakeBench dataset [45] contains 11,000 image-text pairs, specifically designed to evaluate misinformation detection models across a broad range of scenarios. MMFakeBench goes beyond the assumption of single-source forgery and presents samples with “Real”, “Textual Veracity Distortion”, “Visual Veracity Distortion”, and “Cross-modal Consistency Distortion”, making it a more realistic and comprehensive benchmark for assessing multimodal misinformation detection systems. This benchmark is particularly useful in testing the capabilities of models like ours that aim to detect inconsistencies between text and image modalities.
DGM4: DGM4 dataset [61] is a recently released large-scale multimodal manipulation dataset comprising 230,000 image-text paired samples. DGM4 includes over 77,000 pristine pairs and 152,000 manipulated pairs, with manipulations in both image and text modalities. Image manipulation in the dataset involves “face swapping and facial emotion editing”, while text manipulation includes “sentence replacement and textual sentiment editing”. The DGM4 dataset is constructed based on the VisualNews dataset [41], which collects data from multiple news agencies like BBC, The Guardian, USA TODAY, and The Washington Post. This diversity introduces significant distribution discrepancies, making DGM4 ideal for testing the model’s performance under open-world, domain-shift scenarios. The dataset is partitioned into four subsets based on news sources, allowing the evaluation of model robustness in handling different writing styles and regional perspectives.
Table 1: Performance comparison (%) of different models for multi-class classification on the MMFakeBench and DGM4 datasets. The last two highlighted rows represent the results of Lrq-Fact. Bold: best result; Underline: second best result.
Vision-Language Language MMFakeBench (N=1000) DGM4 (N=1000)
[…]
Language Model with 7B Parameter
InstructBLIP [11] Vicuna-7B 7.1 7.9 6.5 7.8 5.3 11.5 4.6 4.5
VILA [39] Llama2-7B 11.5 7.5 25.0 30.0 18.2 27.2 19.0 19.8
BLIP-2 [34] FlanT5-XL 16.4 20.1 27.5 33.0 20.6 32.9 20.7 21.8
LLaVA-1.6 [42] Vicuna-7B 17.4 14.8 25.7 30.8 19.6 30.5 19.0 19.8
Language Model with 13B Parameter
VILA [39] Llama2-13B 11.5 7.5 25.0 30.0 19.4 27.5 19.0 19.8
InstructBLIP [11] Vicuna-13B 13.7 13.2 24.0 28.8 19.2 27.9 19.0 19.8
BLIP-2 [34] FlanT5-XXL 16.7 34.9 27.3 32.8 18.1 26.3 17.9 19.1
LLaVA-1.6 [42] Vicuna-13B 12.0 22.5 25.0 30.0 19.9 29.7 20.1 22.6
Language Model with 34B Parameter
LLaVA-1.6 [42] Nous-Hermes-2-Yi-34B 25.7 44.5 33.7 40.4 32.5 34.3 36.3 39.4
ChatGPT and Llama3.1
GPT-4V-1.7T [53] ChatGPT 51.0 66.8 49.7 54.0 50.6 51.6 55.9 58.5
Paligemma-3B [6] Llama3.1-70B 61.5 67.7 59.1 61.0 54.4 55.1 60.1 60.9
Paligemma-3B [6] ChatGPT 63.3 65.1 61.2 64.3 56.4 57.4 62.1 62.5
4.2 Experiment Details and Settings
For MMFakeBench, we use a sample of 1,000 instances from the validation set, with 300 labeled as real, 300 as textual veracity distortion, 100 as visual veracity distortion, and 300 as cross-modal consistency distortion. A similar sampling approach is applied to the DGM4 dataset, ensuring the same number of samples for each manipulation type. This balanced approach across the different categories ensures a fair comparison between domains.
Our framework integrates both LLMs and VLMs to handle the tasks of question generation and answering. For generating image-focused and text-focused questions, we use the Llama 3.1 [71] model with 70 billion parameters. For each modality, we generate five questions. Rather than fine-tuning Llama 3.1 for this task, we employ in-context learning, where 10 example questions are provided for each modality to guide the model in generating precise, high-quality and relevant questions. To answer the image-focused questions, we leverage Paligemma [6], Google’s vision-language model with 3 billion parameters, which excels at object detection and generating detailed image descriptions. Paligemma’s ability to capture intricate visual contexts makes it particularly effective for visual content analysis. To answer text-focused questions, we employ GPT-4o [3], which is equipped with an extensive knowledge base. This allows the model to provide reliable, well-informed answers, ensuring the accurate verification of textual content. In the decision-making phase, GPT-4o is also used to evaluate the generated question-answer pairs. Known for its strong reasoning and rule-following capabilities, GPT-4o assesses the veracity of the content and offers clear rationales for its conclusions. To further explore the versatility of our framework, we also experiment with using Llama 3.1 70B as both the question-answering model and the decision-maker. All experiments are conducted on NVIDIA 40GB V100 GPUs, providing the computational power required for these extensive evaluations.
4.3 Evaluation Metrics
To assess the performance of the various baselines, we employ a multi-class classification approach. Following the practices established in prior works [86, 56, 9], we use the widely adopted macro-F1 score as our primary evaluation metric. The macro-F1 score provides a balanced measure of precision and recall through their harmonic mean, ensuring fair evaluation across all classes. In addition to the macro-F1 score, we also report macro-precision, macro-recall, and macro-accuracy as complementary metrics, offering a more comprehensive understanding of model performance.
4.4 Comparison Models
We select a diverse range of VLMs with different model sizes as baseline models for comparison. These include: (i) 7B and 13B-parameter models, such as InstructBLIP [11], VILA [39], BLIP-2 [34], and LLaVA-1.6 [42]; (ii) a 34B-parameter model, also represented by LLaVA-1.6; and (iii) the closed-source GPT-4V-1.7T [53] model.
4.5 Experiment Results
We conducted extensive experiments to assess the performance of our proposed framework, Lrq-Fact, in comparison to baseline models on the MMFakeBench and DGM4 datasets.
4.5.1 Main Results.
We provide a comprehensive comparison of various baseline models using the MMFakeBench and DGM4 datasets, as shown in Table 1. Our Lrq-Fact framework employs two different LLMs as decision-makers: GPT-4o, which achieves an F1-score of 63.3%, and Llama-3.1-70B, with an F1-score of 61.5% on MMFakeBench—both outperforming GPT-4V. Notebly, despite having fewer parameters than GPT-4V, our framework consistently delivers better performance, underscoring its efficiency and effectiveness in handling multimodal misinformation detection. Among open-source models, LLaVA-1.6-34B is the leading performer, yet it only achieves an F1-score of 25.7%, significantly trailing GPT-4V, which scores 51%. This reveals a substantial gap in detection capabilities between open-source and proprietary models. Within the same model family, such as LLaVA-1.6-Vicuna-7B and LLaVA-1.6-34B, models with more parameters consistently demonstrate better performance. Smaller VLMs struggle with instruction-following and maintaining high predicted consistency, suggesting that models with 7B parameters lack the multimodal comprehension required to effectively combat misinformation. Due to these limitations in smaller models, we selected mid-sized open-source and proprietary models as baselines for comparison against Lrq-Fact. Further evaluation results for binary classification are provided in the Appendix  A.1
4.5.2 Comparison Against Single-Source Detectors.
We compare Lrq-Fact framework with LLaVA-1.6-34B, GPT-4V-1.7T, and several state-of-the-art single-source detectors [21, 44, 61]. These detectors are representative of the top-performing methods for single-source misinformation detection. Details on each detector are provided in the Appendix  A.2. To ensure a fair comparison, beyond evaluating existing single-source detectors, we also integrate three of the most powerful detectors—FakeingFakeNews [21], LNP [40], and HAMMER [61]—to assess the performance of Lrq-Fact. The results, summarized in Table 2, demonstrate that Lrq-Fact outperforms LLaVA-1.6-34B, GPT-4V, and the single-source detectors across both binary and multiclass classification tasks. This highlights the strong generalization performance of VLMs trained on large-scale corpora in Lrq-Fact, positioning it as a potential baseline framework for future multimodal misinformation detection research.
Table 2: F1-score of single-source detectors on both binary and multi-class misinformation detection tasks for MMFakeBench. TVD, VVD, and CCD represent textual veracity distortion, visual veracity distortion, and cross-modal consistency distortion, respectively. The last highlighted row represents the result of Lrq-Fact.
Model Name Train Source Binary F1$\uparrow$ Multiclass F1$\uparrow$
FakingFakeNews [21] TVD 37.8 -
CNNSpot [76] VVD 23.8 -
UnivFD [52] VVD 28.9 -
LNP [40] VVD 33.0 -
FakeNewsGPT4 [44] CCD 41.7 -
HAMMER [61] CCD 43.0 -
LLaVA-1.6-34B [42] MMFakeBench (N=1000) 62.9 25.7
GPT-4V-1.7T [53] MMFakeBench (N=1000) 72.3 51.0
Paligemma-3B [6] MMFakeBench (N=1000) 76.2 63.3
Figure 3: A case study in multimodal misinformation with cross-modal distortion.
4.5.3 Comparison of Different Categories.
We compare the F1 scores of different VLMs across various misinformation types, including Real, Textual Veracity Distortion (TVD), Visual Veracity Distortion (VVD), and Cross-modal Consistency Distortion (CCD), as shown in Table 3. Across all categories, most open-source models significantly underperform compared to Lrq-Fact and GPT-4V, with the largest gap observed in the handling of textual veracity distortion. This highlights the particular difficulty that open-source models face in accurately detecting textual misinformation.
Table 3: F1-score of different VLMs across various misinformation categories: Real, TVD, VVD, and CCD on MMFakeBench. The last highlighted row represents the result of Lrq-Fact.
[…]
VILA-13B [39] 32.4 13.4 4.3 37.6 21.9
InstructBLIP-13B [11] 41.9 18.8 19.6 23.8 26.0
BLIP2-FLAN-T5-XXL [34] 41.5 39.2 13.1 32.6 31.6
LLaVA-1.6-34B [42] 51.1 37.6 61.7 49.2 49.9
GPT-4V-1.7T [53] 65.3 67.2 57.3 56.5 61.6
Paligemma-3B [6] 66.8 69.2 56.1 66.2 63.3
Case Study. To demonstrate the effectiveness of our method more directly, we showcase an example where our method successfully detects and classifies the multimodal misinformation in Figure 3. More examples can be found in Appendix (see Figure 4, 5, and 6).
5 Conclusion and Future Work
This paper introduces a novel framework, Lrq-Fact, which effectively leverages LLMs and VLMs to enhance multimodal fact-checking. By generating comprehensive, multi-aspect questions for both image and text content, the framework addresses the challenges of detecting misinformation across different modalities. Extensive experiments conducted on two benchmark datasets, MMFakeBench and DGM4, demonstrate that Lrq-Fact significantly outperforms existing methods in terms of detection accuracy and rationale quality. Through its rule-based decision-making module, the framework automates the fact-checking process and provides clear, transparent justifications for its final decision.
Despite these promising results, several directions for future work remain open. First, incorporating external resources, such as Google Search, to retrieve up-to-date information could further enhance the fact-checking process [2, 65]. Second, expanding the analysis of images to include more thorough investigations, like verifying the authenticity of famous individuals depicted, would improve detection capabilities. Third, exploring alternative approaches to ensure the factuality of generated answers, including checking the answers themselves [74, 17], could strengthen the framework’s reliability. Moreover, explorations on the robustness [30, 77, 78, 67, 13] of the system is also important.
A Appendix
A.1 Evaluation Results for Binary Classification.
In addition to multi-class classification, we also assess the binary classification performance to evaluate the overall detection capability of the baseline models across the datasets. Based on the four categories used in the multimodal misinformation detection task, we map these into binary labels. Specifically, instances labeled as “textual veracity distortion”, “visual veracity distortion”, and “crossmodal consistency distortion” are grouped under the "Fake" category, while the label "True" is used to represent real data. Similar to the multi-class evaluation, we employ the widely used F1 score as the primary metric. In addition, precision, recall, and accuracy are reported as supplementary metrics.
Table 4: Performance comparison (%) of different models for binary classification on the MMFakeBench and DGM4 datasets. The last two highlighted rows represent the results from our proposed framework, Lrq-Fact.
Model Language MMFakeBench (1000) DGM4 (1000)
[…]
Language Model with 7B Parameter
InstructBLIP [11] Vicuna-7B 14.7 30.8 13.2 8.1 12.8 33.1 8.2 8.9
VILA [39] Llama2-7B 41.2 35.0 50.0 70.0 26.0 38.2 52.1 67.4
BLIP-2 [34] FlanT5-XL 41.2 35.0 50.0 70.0 25.9 38.3 53.2 68.1
LLaVA-1.6 [42] Vicuna-7B 48.1 48.2 48.5 59.5 37.2 41.2 47.5 54.3
Language Model with 13B Parameter
VILA [39] Llama2-13B 41.1 35.0 50.0 70.0 43.1 32.0 49.2 67.4
InstructBLIP [11] Vicuna-13B 41.1 35.0 49.9 69.9 34.3 48.2 45.7 64.9
BLIP-2 [34] FlanT5-XXL 31.6 63.4 53.6 35.5 30.2 47.2 33.3 31.8
LLaVA-1.6 [42] Vicuna-13B 41.1 35.0 50.0 69.7 34.1 31.2 52.6 65.2
Language Model with 34B Parameter
LLaVA-1.6 [42] Nous-Hermes-2-Yi-34B 62.9 67.1 70.0 63.4 57.1 63.9 67.2 59.3
ChatGPT and Llama3.1
GPT-4V-1.7T [53] ChatGPT 72.3 72.1 72.8 75.6 72.1 74.6 71.8 72.1
Paligemma-3B [6] Llama3.1-70B 72.8 72.1 74.1 75.4 70.9 71.2 72.8 73.1
Paligemma-3B [6] ChatGPT 74.3 76.2 73.1 77.9 72.1 74.3 72.3 74.8
The detailed results are presented in Table 4. Our findings highlight that current models, including both open-source models and GPT-4V, face significant challenges in binary classification tasks on these datasets. Even the advanced GPT-4V model achieves only a 72.3% F1 score with the standard prompt. In contrast, our proposed framework, Lrq-Fact, demonstrates substantial improvements. For example, the BLIP2-FlanT5-XXL model shows a remarkable 1% increase in F1 score when integrated into our framework. This improvement can be attributed to the effective incorporation of rationales, actions, and tool usage, which enhances multimodal understanding in the context of misinformation detection.
A.2 Details of Existing Single-Source Detectors
In this section, we provide detailed descriptions of the single-source detectors used as baselines in our experiments. These detectors represent state-of-the-art models for detecting misinformation in specific modalities, including textual and visual content.
FakeingFakeNews.
FakeingFakeNews [21] is designed to detect textual fake news, focusing particularly on human-written misinformation. It introduces an innovative method for generating training data, leveraging well-known techniques used in human-authored propaganda. The framework uses the ROBERTA [46] model as its backbone and is trained on the PROPANEWS dataset. In our experiments, we applied the ROBERTA detector with its default settings, as provided in the FakeingFakeNews framework.
CNNSpot.
CNNSpot [76] specializes in detecting artificial images generated by models. It utilizes the ResNet-50 model as its classifier backbone. CNNSpot highlights the benefits of data augmentation, such as JPEG compression and Gaussian blur, which improve the detector’s generalization capabilities. In our study, we used the pre-trained CNNSpot model with default hyperparameters to detect visual veracity distortions.
UnivFD.
UnivFD [52] is a versatile fake image detector that operates in a feature space not explicitly trained to differentiate between real and fake images. By accessing the feature space of a pre-trained vision-language model, UnivFD applies a nearest-neighbor approach to detect fake images from various sources. The use of a large pre-trained model enables smoother decision boundaries, enhancing generalization. We utilized the pre-trained UnivFD model with default settings to perform visual veracity distortion detection in our experiments.
LNP.
LNP [40] employs a denoising model to extract noise patterns from spatial images, then analyzes the frequency domain of these noise patterns to detect fake images. It uses the ResNet-50 model as its backbone for classification. In our study, we used the pre-trained LNP detector with default hyperparameters to carry out visual veracity distortion detection.
HAMMER.
HAMMER [61] is a multimodal detector developed to identify multimedia manipulation. Built on the pre-trained vision-language model ALBEF [35], it uses two unimodal encoders and a multimodal aggregator. HAMMER employs hierarchical manipulation reasoning: shallow reasoning checks semantic alignment between image and text embeddings, while deep reasoning performs cross-modal fusion for manipulation detection. In this work, we used the off-the-shelf HAMMER detector with default hyperparameters to detect cross-modal consistency distortion.
FakeNewsGPT4.
FakeNewsGPT4 [44] is based on large vision-language models (LVLMs) and is designed to detect multimodal fake news. It integrates two types of forgery-specific knowledge: semantic correlation and artifact tracking. The model enhances LVLMs by incorporating these two elements, utilizing a multi-level cross-modal reasoning module for semantic correlation and a dual-branch fine-grained verification model for unimodal localized detail comprehension. In our experiments, we used the pre-trained FakeNewsGPT4 with default hyperparameters to detect cross-modal consistency distortion.
A.3 Example Analysis
Real. In this case study (Figure 4), the description is well-constructed and aligns perfectly with the image and the textual context. The depiction of Jake Davis as a young man in casual clothing, standing in a relaxed manner, accurately reflects the narrative of his release from a young offender institution. The visual context provided in the image adds credibility to the news article, confirming the validity of the description. There are no discrepancies between the image and the text, making the description not only good but also a reliable tool to confirm the factual correctness of the news.
The questions presented in this case are well-formed and reliable. They are designed to extract key details from both the image and the text, ensuring comprehensive verification. The image-focused questions effectively ask about the setting and identity, which helps in confirming whether the person and location in the image match the article’s claims. The text-based questions aim to validate the timeline and factual details, ensuring a consistent narrative. These questions are precise and structured to get the best possible answers, making them a solid mechanism for cross-verifying facts.
Textual Veracity Distortion. The description in Figure 5 accurately depicts a lighthouse in a Gothic architectural style, positioned on a rocky shore with surrounding water and seagulls. The image is valid and corresponds with the article’s general theme. However, the description’s alignment with the actual claim in the text—that the lighthouse is haunted and located in Greece—proves to be incorrect. While the description is visually consistent and good, it does not support the erroneous textual claim, showing how important it is to assess both text and visuals in tandem.
The questions in this case are reliable and appropriately structured to identify discrepancies between the image and the text. The image-focused questions ask about the architectural style and contextual clues from the image, while the text-based questions explore the factual accuracy of the claim that this lighthouse is haunted and located in Greece. The questions provide a good framework for fact-checking by encouraging thorough scrutiny of both visual and textual elements. This ensures that any distortions or misrepresentations in the article are effectively highlighted, making the questions a valuable tool for getting to the truth.
Visual Veracity Distortion. In Figure 6, the description of a clock tower in yellow and white is valid and clear, but the image itself shows a structure that is clearly gold and digitally altered. The description is good in terms of clarity and helping readers visualize the article’s claim, even though it does not reflect the manipulated nature of the image. This highlights the importance of analyzing the veracity of visuals alongside textual descriptions.
The questions are well-crafted to reveal any visual inconsistencies. The image-focused questions ask about the color and reality of the clock tower, which are key to identifying that the clock tower has been digitally altered. The text-based questions, which probe the existence of such a clock tower in real life, also help uncover discrepancies. These questions are reliable and precise, aimed at extracting the best possible answers and guiding the evaluation of the article’s claims against the evidence provided by the image.
Cross-modal Consistency Distortion. This case (Figure 7) involves a clear mismatch between the text and the image, where the article describes a little girl holding uncooked rolls, while the image shows her holding paper towels. The description is coherent and well-explained, making it a good tool to visualize the scenario presented in the article. However, the inconsistency between the image and the text highlights a cross-modal distortion. Despite this, the description itself remains valid in its own right.
The questions presented are well-designed to highlight the inconsistency between the text and the image. The image-focused questions ask about the scene and the object the girl is holding, providing clear answers that reveal the mismatch. The text-based questions further confirm this by addressing the article’s lack of accurate description. These questions are well-structured and reliable, allowing for an in-depth examination of both the image and the text to expose cross-modal discrepancies. They guide the analysis toward the best possible answers by focusing on the key elements that need verification.
A.4 Instruct Prompt for Lrq-Fact
The Lrq-Fact framework employs a series of structured prompts to guide LLMs and VLMs in multimodal fact-checking. These prompts facilitate the generation of detailed image descriptions, contextually relevant questions, and well-informed answers that probe the veracity of both visual and textual content. In the final step, a rule-based decision-maker evaluates the generated questions and answers to provide a final judgment on the consistency between the text and image, ensuring accurate detection of misinformation.
Image Description Prompt. The first step is to generate a detailed description of the image, capturing all relevant elements that help assess its consistency with the textual content. This description is crucial for identifying potential inconsistencies or manipulations between the image and the accompanying article. The specific prompt used to generate this description is provided in Figure 8.
Image-Focused Questions Prompt. This stage generates relevant image-focused questions designed to verify the accuracy, authenticity, and relevance of the visual content in relation to the article. These questions help clarify the image content and assess its relation to the text. The specific prompt for generating these questions is illustrated in Figure 9.
Image-Focused Answers Prompt. After generating the image-focused questions, this prompt helps in generating answers that analyze the visual content directly from the image. These answers are based on the key elements and actions identified in the image, ensuring that the responses are relevant and insightful. The specific prompt for this is shown in 10.
Text-Focused Questions Prompt. To critically assess the factual claims in the text, this prompt generates relevant questions targeting specific elements such as dates, names, locations, and events. The generated questions aim to challenge the accuracy of the claims made in the article. The specific prompt used for text-focused questions is shown in Figure 11.
Text-Focused Answers Prompt. After generating the text-focused questions, this prompt helps generate accurate answers using the model’s built-in knowledge. These answers verify factual accuracy and challenge any unsupported claims in the article. The specific prompt for this is illustrated in Figure 12.
Rule-Based Decision-Maker Prompt. After gathering information from the image and text analyses, the rule-based decision-maker evaluates the consistency between modalities and makes a final determination about the article’s veracity. This module provides a detailed explanation for the final judgment. The specific prompt for the rule-based decision-making process is shown in Figure 13.
Figure 4: This example case illustrating the alignment between image and text in a fact-checking process. The generated questions verify key elements, ensuring consistency and accuracy in multimodal misinformation detection. This demonstrates how targeted questions and well-constructed descriptions enhance reliable fact-checking outcomes.
Figure 5: Example case illustrating textual veracity distortion. The image description aligns visually with the content, but fails to support the false textual claim about the haunted lighthouse’s location in Greece. The generated questions are designed to detect inconsistencies, providing a thorough framework for fact-checking by scrutinizing both visual and textual elements.
Figure 6: This example case highlighting visual manipulation. The description accurately conveys the textual claim about a yellow and white clock tower, but fails to reflect the digitally altered gold structure seen in the image. The questions focus on detecting visual discrepancies, such as the altered colors, and also probe the existence of such a clock tower, providing a reliable framework for evaluating both the image and text.
Figure 7: This example case demonstrating cross-modal distortion. The description is clear and helps visualize the article’s scenario of a little girl holding uncooked rolls, while the image actually shows her holding paper towels. This mismatch between text and image points to a cross-modal distortion. The questions are well-crafted to reveal this inconsistency by focusing on both the scene and the object in the girl’s hands, providing a reliable framework for identifying the discrepancy between the article and the image.
Image Description Prompt: Please provide a detailed and comprehensive description of the image shown. Focus on identifying all visible elements including objects, people, setting, and any interactions or actions taking place. Describe the colors, textures, mood, and any other notable aspects that contribute to the overall context and significance of the image.
Figure 8: Structured prompt to generate detailed image descriptions.
Image-Focused Questions Prompt: Given the following news article [news text], generate up to [number of questions] questions that are relevant to the news article and focused on elements visible in the image. These questions should help clarify the image content, assess its authenticity, and explore its relation to the news article.
Instructions for Question Generation:
Focus on generating high-quality questions that are directly relevant to both the news article and the visual elements in the image. The questions should investigate visible interactions, settings, actions, any text or symbols, and specific individuals or objects mentioned in the article. Additionally, include questions that assess the authenticity of the image, such as whether the image could have been generated by AI, or whether there are any unusual or suspicious elements in the image.
Aim to Generate:
- The most important questions that will help determine whether the image supports or contradicts the context provided by the news article.
- Questions that may uncover discrepancies or inconsistencies between the image and the news article.
- Questions that assess the authenticity or realism of the image, particularly in relation to the article’s content.
Avoid asking for:
- Personal details not visible in the image.
- Information that cannot be inferred from the image itself.
Example Questions:
1. What event is depicted in this image?
2. How are the people in the image interacting?
3. Is the person in the image performing [action from article]?
4. What are the technical aspects or tools used to create this image?
5. What emotions does this image evoke?
Questions: 1. , 2. , …
Figure 9: Structured prompt to generate relevant image-focused questions.
Image-Focused Answers Prompt: You are an advanced AI model with access to a vast repository of knowledge and the capability to analyze both text [news text] and image [image]. Ensure your answer is directly related to the question [generated questions] and it is based on the image. Answers: 1. , 2. , …
Figure 10: Structured prompt to generate answers for the image-focused questions.
Text-Focused Questions Prompt: Given the following news article [news text], analyze the text and formulate up to [number of questions] questions that probe the accuracy and verifiability of the information contained in the article. These questions should be designed to identify potential inaccuracies or areas that can be confirmed or challenged based on general knowledge or the text itself.
Instructions for Question Generation:
Focus on generating high-quality, fact-checking questions that can be answered directly through general knowledge that an LLM might possess. Identify and question significant factual claims, examine dates, locations, names, and other data mentioned in the article, and challenge any assumptions. The goal is to produce questions that facilitate direct verification of the facts stated in the article.
Aim to Generate:
- Questions that challenge the accuracy of specific claims made in the article and can be answered based on general knowledge.
- Questions that explore potential inconsistencies or contradictions within the article’s content.
- Questions that assess the logical coherence and factual basis of the article’s claims.
Avoid asking for:
- Information requiring external sources or verification beyond general knowledge.
- Speculative or opinion-based questions.
Example Questions:
1. Does the description of the “meeting between world leaders on March 5th” align with the known schedule of diplomatic events for that time?
2. Is the account of “a large protest taking place in front of City Hall” consistent with known reports of protests in that area during the stated period?
3. Does the timeline of “economic sanctions being imposed after the incident” logically follow the typical process for such actions?
4. Are the historical events referenced, such as “the financial crisis of 2008”, accurately portrayed in the article?
Questions: 1. , 2. , …
Figure 11: Structured prompt to generate relevant text-focused questions.
Text-Focused Answers Prompt: You are an advanced AI model with access to a vast repository of knowledge. Your task is to answer the following questions [generated questions] based on your built-in knowledge. While a news article [news text] is provided for context, you must answer the questions solely based on your own knowledge and not refer to the article’s content.
Instructions for Answering:
Provide accurate, clear, and concise answers to each question. Your responses should be based entirely on your general knowledge and the information you have learned. Do not reference or rely on the content of the provided news article when forming your answers. Each answer should be factually correct and directly relevant to the question asked.
Answers: 1. , 2. , …
Figure 12: Structured prompt to generate answers for the relevant text-focused questions.
Rule-Based Decision-Maker Prompt: Your objective is to determine whether the article and image are real or fake by analyzing the following information: 1. News Article Text: [news text]
2. Image Description: [image description]
Note: This description helps verify consistency with the news text. It is generally reliable but may contain minor discrepancies, such as using different terms like “ocean” instead of “water”.
3. Generated Image-Focused Questions and Answers: [generated image-focused QAs]
Note: These answers were generated by an AI and may contain mistakes, such as incorrect details regarding locations, names, dates, or objects. They might also incorrectly suggest that the image has been manipulated or is AI-generated. If the answers suggest manipulation or that the image is AI-generated, this should have very low effect on your final decision, especially if the image description and news article text do not contain such indications.
4. Generated Text-Focused Questions and Answers: [generated text-focused QAs]
Note: These are based on the knowledge of GPT4-O, which is generally reliable but prone to hallucinations or contradictions with other provided information.
Instructions:
To make an accurate judgment of the multimodal misinformation, please follow these steps:
Step 1. Is there any credible objective evidence refuting the news description? If yes, assign the label: Textual Veracity Distortion. If no, continue to Step 2.
Step 2. Is there any credible objective evidence refuting the news image? If yes, assign the label: Visual Veracity Distortion. If no, continue to Step 3.
Step 3. Does the news caption match the content of the news image? If no, assign the label: Mismatch. If yes, and none of the above applies, assign the label: Real.
Additional Guidelines:
1. Assess Overall Consistency: …
2. Examine Details: …
3. Analyze Facial Expressions and Body Language: …
4. Identify Unrealistic Elements: …
5. Cross-Modal Consistency: …
6. Final Judgment: …
7. Select the Most Relevant Label: …
8. Provide a Detailed Explanation: …
Example Output:
1. Final Judgment: Fake
2. Label: Visual Veracity Distortion
3. Explanation: The image description mentions a “cat with pink eyes”, which is highly unnatural and suggests the image is AI-generated. Additionally, ….
1. Final Judgment: [Real or Fake]
2. Label: [Select one: Textual Veracity Distortion, Visual Veracity Distortion, Mismatch, Real]
3. Explanation: [Provide your explanation here]
Figure 13: Structured prompt to make the final decisions and provide an explanation.